what is the slope of a line whose graph is parallel to the graph of -5x+y=7? perpendicular to the graph 3x+y=7?
Answer by stanbon(75887)

what is the slope of a line whose
graph is parallel to the graph of -5x+y=7?
y = 5x+7
slope = 5
The slope of every line parallel to it will have slope = 5
-----------------------------------------------------------
perpendicular to the graph 3x+y=7?
y = -3x+7
slope = -3
The slope of ever line perpendicular to it will have slope = 1/3
==================
Cheers,
Stan H.
==============